OpenStudy (czarluc):

Sin (4x) + sin(2x) =0

OpenStudy (anonymous):

hint: sin(4x) = sin(2*2x) = 2sin(2x)cos(2x)

OpenStudy (czarluc):

Nope. Its spsssin 4x +sin2x=0

OpenStudy (czarluc):

Sourwing, I know that but only there :(

OpenStudy (czarluc):

Oh OK thanks

OpenStudy (solomonzelman):

looks confusing where the powers and where the sin of (2x) and (4x)....

OpenStudy (czarluc):

I changed it already @SolomonZelman

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \sin (4x) + \sin(2x)=0 }\) \(\LARGE\color{black}{ 2\sin (2x)\cos(2x)+ \sin(2x)=0 }\) \(\LARGE\color{black}{ (2\cos(2x)+1)\times \sin(2x)=0 }\)

OpenStudy (solomonzelman):

(zero product property)

OpenStudy (czarluc):

Sorry, but i don't get it for the third line

OpenStudy (solomonzelman):

I factored out of `sin(2x)`

OpenStudy (czarluc):

Like, how did sin(4x) become2cos(2x)+1

OpenStudy (solomonzelman):

sin(4x) became 2sin(2x)cos(2x) using the sin(a+b) rule.

OpenStudy (anonymous):

look how can you be at two things at once

OpenStudy (solomonzelman):

??

OpenStudy (anonymous):

look you hacker

OpenStudy (czarluc):

But, isn't it it should become 4 cos^3 x sinx- 4sin^3 x cosx?

OpenStudy (anonymous):

Solomon

OpenStudy (solomonzelman):

would would sin(4x) become that?

OpenStudy (czarluc):

Yeah,by using the gen. Add formula

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \sin(4x)=\sin(2x+2x) }\) \(\LARGE\color{black}{ \sin(4x)=\sin(2x)\cos(2x)+\sin(2x)\cos(2x) }\) \(\LARGE\color{black}{ \sin(4x)=2\sin(2x)\cos(2x) }\) \(\LARGE\color{black}{ \color{red}{ \sin(4x)}+\sin(2x)=0 }\) \(\LARGE\color{black}{ \color{red}{ 2\sin(2x)\cos(2x)}+\sin(2x)=0 }\) Good?

OpenStudy (czarluc):

Yeah, and I can't seem to understand what to do after that

OpenStudy (czarluc):

Thanks btw

OpenStudy (czarluc):

Oh nvm, I understood it already, thanks Solomon!

OpenStudy (solomonzelman):

Okay, and then you get, \(\LARGE\color{black}{ 2\color{blue}{\sin(2x)}\cos(2x)+\color{blue}{\sin(2x)}=0 }\)

OpenStudy (solomonzelman):

Factor it out of sin(2x). What do you get?

OpenStudy (czarluc):

It will become (2cos (2x)+1)(sin(2x)=0

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

And then applying the zero product property, you get `sin(2x)=0` or `2cos(2x)+1=0`

OpenStudy (solomonzelman):

can you solve the sin(2x)=0 for x?

OpenStudy (czarluc):

And now my problem is how to get the values for x

OpenStudy (czarluc):

Sin (2x)=0 is x=0,180& 360 right?

OpenStudy (czarluc):

oops I mean, x= 0,90,and 180?

OpenStudy (solomonzelman):

yes, x=90 and 180. so that the sin(2x) is sin(180) sin(360) as you have just fixed it... :)

OpenStudy (solomonzelman):

and the second part, `2cos(2x)+1=0` subtract 1 from both sides, divide both sides by 2 then tell me cos(a) for which value(s) of a is 1/2?

OpenStudy (czarluc):

a= 60 and 120?

OpenStudy (solomonzelman):

sure that cos(120) is 1/2?

OpenStudy (solomonzelman):

I wouldn't say that.. lol

OpenStudy (czarluc):

isn't it correct? as when 2a= 120,240,480 and 600 and a now is 60,120,240 and 300?

OpenStudy (czarluc):

nvm 480,600 and 240,300

OpenStudy (solomonzelman):

I mean sin(2x)=1/2 is when x=30.

OpenStudy (czarluc):

but isn't it that we are using the cos function? and cos(2x)= -1/2

OpenStudy (solomonzelman):

yeah, my bad cos.

OpenStudy (solomonzelman):

cos(60)=1/2 so when 60 is 2x, then x is 30.

OpenStudy (solomonzelman):

It was my bad that I typed sin when I mean cos -:(

OpenStudy (czarluc):

it's ok, but is the values right? i mean 60 and 120 for a?

OpenStudy (solomonzelman):

HINT: \(\large\color{black}{ 30 }\) degrees is \(\large\color{black}{ \pi/6 }\) .

OpenStudy (solomonzelman):

we are doing the 2x, so 2x would be \(\large\color{black}{ \pi/3 }\) (when x is \(\large\color{black}{ \pi/6 }\) )

OpenStudy (czarluc):

but it is negative. so isn't it that 2x would be 120 and 240? because it is where cos is negative? :3

OpenStudy (solomonzelman):

yes.

OpenStudy (czarluc):

so it is 60 and 120?

OpenStudy (solomonzelman):

yees

OpenStudy (czarluc):

oh ok, thanks @SolomonZelman !!!

OpenStudy (solomonzelman):

No problem.